Can I please get help with this problem.
x/x-4 and g(x)=-7/x and solve for (gof)(x)
Thank you so much

Answer by rothauserc(4718)

f(x) = (x / (x-4))
g(x) = -7x
:
gof means apply f first and then g, then
(gof)(x) = -7(x / (x-4)) = -7x / (x-4)